write the equasion of a circle with a diameter having endpoints at
(-2,6) and (8,4)?
: write the equasion of a circle with a diameter having endpoints at
(-2,6) and (8,4)?

Answer by scott8148(2761)

the center is midway between the endpoints at (3,5)

the radius squared is the square of the distance from the center to an endpoint
__ r^2=5^2+1^1

(x-3)^2+(y-5)^2=26